Use implicit differentiation to obtain dy over dx

Cosx +sinx=xsquared+ysquared


1
Expert's answer
2021-05-20T08:37:20-0400

Ans:-

"Cosx +Sinx=x^2+y^2\\" then "y=\sqrt{Cosx+Sinx-x^2}"

Differentiate both side with respect to x

"\Rightarrow -Sinx+Cosx=2x+2y\dfrac{dy}{dx}\\
\Rightarrow \dfrac{dy}{dx}=\dfrac{Cosx-Sinx}{2x+2y}\\"


"\Rightarrow \dfrac{dy}{dx}=\dfrac{Cosx-Sinx}{2(x+\sqrt{Cosx+Sinx-x^2})}"